what is the kinetic energy of Kenny, a 65-kg runner who is running at 10m/s


1
Expert's answer
2021-02-15T17:03:52-0500
"KE=\dfrac{1}{2}mv^2,""KE=\dfrac{1}{2}\cdot65\ kg\cdot(10\ \dfrac{m}{s})^2=3250\ J."